A population consists of the values (1,3,5,7) consider a sample size of two that can be drawn from this population

1
Expert's answer
2022-02-28T16:10:07-0500

We have population values "1,3,5,7" population size "N=4" and sample size "n=2."

Thus, the number of possible samples which can be drawn without replacement is "\dbinom{4}{2}=6."


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Sample\ values & Sample\ mean(\bar{X}) \\ \hline
 1,3 & 2\\
 \hdashline
 1,5 & 3\\
 \hdashline
 1,7 & 4\\
 \hdashline
 3,5 & 4\\
 \hdashline
 3.7 & 5\\
 \hdashline
 5,7 & 6\\
 \hdashline

\end{array}"



The sampling distribution of the sample mean "\bar{X}" is


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 & \bar{X} & f & f(\bar{X}) & Xf(\bar{X})& X^2f(\bar{X}) \\ \hline
 & 2 & 1 & 1/6 & 2/6 & 4/6\\
 \hdashline
 & 3 & 1 & 1/6 & 3/6 & 9/6 \\
 \hdashline
 & 4 & 2 & 2/6 & 8/6 & 32/6\\
 \hdashline
 & 5 & 1 & 1/6 & 5/6 & 25/6 \\
 \hdashline
 & 6 & 1 & 1/16 & 6/6 & 36/6 \\
 \hdashline
 Sum= & & 6 & 1 & 4 & 53/3\\
 \hdashline
\end{array}"


"\mu=\dfrac{1+3+5+7}{4}=4"

"\sigma^2=\dfrac{1}{4}((1-4)^2+(3-4)^2+(5-4)^2+(7-4)^2)"

"=5"


Check

The mean of the sample means is


"\mu_{\bar{X}}=E(\bar{X})=4=\mu"




"Var(\bar{X})=\sigma^2_{\bar{X}}=E(\bar{X}^2)-(E(\bar{X}))^2"

"=\dfrac{53}{3}-(4)^2=\dfrac{5}{3}"

The standard deviation of the sample means is


"\sigma_{\bar{X}}=\sqrt{\sigma^2_{\bar{X}}}=\sqrt{5/3}"

"\dfrac{\sigma^2}{n}\cdot\dfrac{N-n}{N-1}=\dfrac{5}{3}\cdot \dfrac{4-2}{4-1}=\sigma^2_{\bar{X}}"

"\mu_{\bar{X}}=\mu"

"\sigma_{\bar{X}}=\dfrac{\sigma}{\sqrt{n}}\sqrt{\dfrac{N-n}{N-1}}"